OpenStudy (anonymous):

If sinβ is the geometric mean between sinα, cosα then cos2βisequal to

OpenStudy (anonymous):

first sin is sin beta then sin alpha,cos alpha,cos 2beta

Parth (parthkohli):

OK.\[\sin^2 \beta = \sin\alpha \cdot \cos\alpha\]And we know that,\[\cos(2\beta) = 1 - 2\sin^2(\beta)\]

OpenStudy (anonymous):

\[if \sin \beta,is the geo mean btw \sin \alpha,\cos \alpha,then \cos 2\]

OpenStudy (anonymous):

\[\beta 2 \iota [ 3 x 4 ] \]

OpenStudy (anonymous):

@Rahul .cant able to follow

Parth (parthkohli):

Do you understand my solution?

OpenStudy (anonymous):

thats what i have studided

OpenStudy (anonymous):

sin beta=sin alpha cos alpha

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

ParthKohli you are wrong in one thing

Parth (parthkohli):

If \(b\) is the geometric mean between \(a\) and \(c\), it means that \(b^2 = ac\). By the above property, it can be said that \(\sin^2 \beta = \sin\alpha \cdot \cos \alpha\). We need to determine the value of \(\cos(2\beta)\) which is equal to \(1- 2 \sin^2 \beta \). Plug in the value of \(\sin^2 \beta.\)

OpenStudy (anonymous):

\[1-2\sin \alpha \cos\]

Parth (parthkohli):

That's right. You can simplify it further.

OpenStudy (anonymous):

1-2 sin alpha cos alpha

OpenStudy (anonymous):

Parth Kohli you know what

OpenStudy (anonymous):

1-sin 2 alpha

Parth (parthkohli):

That's right!\[\therefore \cos(2\beta) = 1 - \sin(2\alpha)\]

OpenStudy (anonymous):

ParthKohli more like Virat Kohli the cricketer

Parth (parthkohli):

lol

OpenStudy (anonymous):

common ler become fan of each other

OpenStudy (anonymous):

then

OpenStudy (anonymous):

alpha can be divided

OpenStudy (anonymous):

how

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

if you get the divisor then because the dividend is alpha

OpenStudy (anonymous):

how

OpenStudy (anonymous):

you have to get the divisor by its formula

OpenStudy (anonymous):

there must be a formula

OpenStudy (anonymous):

\[If \sin \beta is the geometric mean \between \sin \alpha,\cos \alpha then \cos 2\beta is equal \to\]

Parth (parthkohli):

I've already given you the answer...

OpenStudy (anonymous):

@ParthKohli the problem didnt get finished

OpenStudy (anonymous):

see question 29

OpenStudy (anonymous):

oh u live in india !

OpenStudy (anonymous):

seeing the question

OpenStudy (anonymous):

hmm..lol

OpenStudy (anonymous):

question taken from online exam.i was searching through net

ganeshie8 (ganeshie8):

continuing parth's wrok : \[\large \begin{array}\\ 1-\sin 2\alpha &= \sin^2\alpha + \cos^2\alpha - 2\sin\alpha \cos\alpha \\&= (\sin \alpha - \cos \alpha)^2 \end{array} \]

ganeshie8 (ganeshie8):

\[\large \begin{array}\\ &= \color{red}{2}(\sin \alpha \color{red}{\cos(\pi/4) }- \cos \alpha\color{red}{\sin(\pi/4)})^2 \end{array} \]

ganeshie8 (ganeshie8):

\[\large \begin{array}\\ &= \color{red}{2}(\sin (\alpha-\pi/4) )^2\end{array} \] \[\large \begin{array}\\ &= \color{red}{2}\sin^2 (\pi/4-\alpha )\end{array} \]